Shift summary for day shift 7/12

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Mon, 12 Jul 1999 14:48:46 -0600 (MDT)

Shift Summary for day shift, 12/July/99

On Shift: Ron Ray, Jim Hirschauer

Hours of beam delivered at all intensities: 8.25
Hours of beam delivered at nominal intensity: 8.25
Hours of E832 data (special runs not included): 6.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 1.4 hrs - ktev1,tps2d1 failure, fire alarm and
ensuing evacuation, banana problems

========================================================================
Runs ended this shift:
Run 13861 - 24 million events - BEAM_832
Run 13865 - aborted @ tape drive failure - BEAM_832
Run 13867 - in progress - BEAM_832

Tapes brought to FCC this shift: QKC119,QKC120,QKC121,QKC122,QKC123,QKC124
QKS119,QKS120,QKS121,QKS122,QKS123,QKS124
QKK119,QKK120,QKK121,QKK122,QKK123,QKK124
QKB041,QKB042
plus various onsplit tapes included in
the soon to follow adendum

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Not-so-smooth 832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13861 7:00 take over from day shift, run in progress

7:58 ktevana errors, ran kfrend

9:03 beam lost for ~10 minutes. Marj cycles the
voltage on DC3, because the current draw has been
creeping up. DC group will take a look at this
during access on Tues. Shiftees should keep an
eye on the DMM in front of DC current monitors, if
it goes above 5 uA call Marj.

9:15 beam returns, DC3Y off spill current is 3 uA @
2200V. Cycling HV made no difference

11:15 We miss a few spills and tapes are nearly full, so
we stop run 13861 and change tapes. Problems are
encountered setting the DA to recognize the
mounted tapes. Further complications when ktev1,
tps3d1's unload button gets stuck. We do a fresh
start.

13865 11:40 Start run 13865.

11:50 Run 13865 aborts.
ktev1, tps2d1 kicks. Since run had only a few
spills we move to tps3d2 and overwrite tapes.
Run 13865 will not appear on tapes.

12:00 Fire alarm goes off. We leave the building, fire
dept shows up. KTeV hall is reduced to ashes in
towering inferno. I mean, it seems to be a false
alarm generated by CO2 panel.

12:30 Fresh start after tape drive reconfiguration.
Problem with bananas, can't unhang triggers.

13867 13:00 Pat "Banana-Man" Toale manually reloads maps.
After a few kfrends we are back in business.
Start Run 13867.

13:39 Slow Daq gives warning:
Crate 10, Rack 165 -> DTM front panel switch is on
"tracking." Theo A is called and says that this
happens occasionally and should go away. He is
correct.

3:30 Hand over to evening shift. Run 13867 in progress

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o Do Laser_Scan at anticipated downtime (if enough time), do Muon_Run and
magnet flip upon beam recovery.